The heat generated in a circuit is given by H = I2Rt joule where I is current, R is resistance and t is time. If the percentage errors in measuring I, R and t are 1%, 2% and 1% respectively, the maximum error in measuring heat will be

Grade:Upto college level

1 Answers

Saurabh Singh
10 years ago
( error in H) /H= 2*(error in I)/I +( error in R)/R +( error in t)/t

%error in H = 2*1+2+1
=5%

Thanks & Regards


Saurabh Singh,


askIITians Faculty


B.Tech.


IIT Kanpur